# “Add more config detail”: A Taxonomy of Installation Instruction Changes

### v1

@@ Abstract
README files play an important role in providing installation-related instructions to software users and are widely used in open source software systems on platforms such as GitHub. However, these files often suffer from various documentation issues, leading to challenges in comprehension and potential errors in content. Despite their significance, there is a lack of systematic understanding regarding the documentation efforts invested in README files, especially in the context of installation-related instructions, which are crucial for users to start with a software project. To fill the research gap, we conducted a qualitative study, investigating 400 GitHub repositories with 1,163 README commits that focused on updates in installation-related sections. Our research revealed six major categories of changes in the README commits, namely pre-installation instructions, installation instructions, post-installation instructions, help information updates, document presentation, and external resource management. We further provide detailed insights into modification behaviours and offer examples of these updates. Based on our findings, we provide recommendations to practitioners for maintaining their README files, as well as motivations for future research directions. These recommendations and research directions encompass completeness, correctness and up-to-dateness, and information presentation consideration. The proposed research directions span the development of automated documentation tools and empirical studies to enhance comprehension of the needs of documentation users. Furthermore, we provide a comprehensive README template tailored to cover the installation-related sections for document maintainers, serving as a practical starting point for their efforts.
@@
Software documentation, README Files, Qualitative Analysis, Thematic Analysis
1 Introduction
@@
Good software documentation serves as a vital source of information for different stakeholders involved in software systems. In particular, README files are among the first software documents that users, such as other developers and end-users, and newcomers encounter when accessing an Open Source Software (OSS) project, leaving them with a crucial first impression and providing them with an introductory overview of the project [1, 2]. At the time of this study, GitHub, the most popular open source code platform, hosts more than 372 million repositories, [^1] http://web.archive.org/web/20230424004513/https://en.wikipedia.org/wiki/GitHub underscoring the high demand for maintaining high quality README files.
@@
One of the most common types of content in GitHub README files is related to basic instructions (e.g., project installation guide) [1]. A well-written “How” section could alleviate the effort for developers and users, especially newcomers, to get hands-on with the projects. In fact, overcoming the challenge of setting up an environment is one of the hurdles for newcomers to contribute to OSS projects [3].
@@
However, software tutorials that focus on installing software tools, configuring virtual computing environments, and other related continuous deployment processes have been found to not be executable in more than half of the cases investigated by related work [4]. Furthermore, previous empirical studies identified inappropriate installation instructions, outdated content, and content clarity as frequently appearing documentation issues [5, 6]. These factors collectively constitute the impediments for both developers and project users to effectively use the software repositories.
@@
Although previous studies have pinpointed software documentation issues, particularly in installation-related topics [5, 6, 7], their analyses were broad in scope, with the aim of synthesising different aspects of documentation issues from surveys or software documents. In contrast, documentation updates provide a new perspective consisting of the modifications and two states: before and after the modification. In fact, increased documentation activities and efforts have demonstrated their efficacy in improving the popularity of and encouraging the participation in GitHub projects [8]. Considering the documentation issues and the benefits of investing efforts to maintain high-quality documentation, knowledge of the practices of the README file efforts becomes essential. This knowledge not only can provide practitioners with recommendations, but can also provide researchers with insight for the development of automated documentation tools. However, empirical evidence to support practices related to software documentation efforts remains scarce.
@@
In this work, we conducted an empirical study to qualitatively analyse README file modification activities in installation-related topics to fill the research gap. We mined GitHub repositories from GHArchive [^2] https://www.gharchive.org/ that were active between 2015-01-01 and 2023-02-28, applying stringent filters on different features for the repositories, resulting in 12,908 projects. We examined a statistically significant sample of 1,168 README file commits within 400 randomly selected repositories and adopted thematic analysis [9] to synthesise the types of README update tasks in the modified sections.
@@
Based on our analysis, we synthesised a comprehensive taxonomy of 189 update behaviours, which are related to the categories of: (1) pre-installation instructions, (2) installation instructions, (3) post-installation instructions, (4) help information, (5) document presentation, and (6) external resource management. We further discussed their implications for documentation maintainers and software engineering researchers based on our discoveries from perspectives of completeness, correctness and up-to-dateness, and information presentation considerations. In particular, we provided a README file template tailored to cover the installation-related sections for document maintainers to get start with.
Our key contributions are as the followings:
  • A large-scale repository dataset, which can be used for future research.
  • Formalisation of a detailed list of software development tasks related to README file content.
  • A comprehensive taxonomy for the types of installation-related documentation modifications for OSS projects.
  • Actionable recommendations for OSS documentation maintainers and researchers.
2 Related Work
2.1 Empirical Studies on Software Documentation
@@
One source of information for researchers is directly from software engineering practitioners. Several studies have investigated the impact of software documentation. Kajiko-Mattson [10] studied the practices of software documentation in Swedish organisations, identifying rudimentary documentation requirements relevant within corrective maintenance. Garousi et al. [11] conducted an industrial case study on the usage and usefulness of different documentation artefacts in the software development life cycle. Their findings indicated variations in documentation usage based on information needs, and the recommendations derived from their study have been integrated into the company’s practices where the experiment took place. Software documentation issues are also investigated. Uddin and Robillard [12] conducted a large-scale survey on API documentation failure factors among IBM practitioners, and identified the most pressing issue related to content, as opposed to presentation. Aghajani et al. [6] presented the views of practitioners on various issues of software documentation, along with their proposed solutions to address these issues. Moreover, the quality of software documents is another perspective that has been studied. Treude et al. [13] recruited technical editors and synthesised a framework to assess the quality of the software documentation from the respondents. Their work provides visions for a potential unified quality framework for assessing software documentation.
@@
However, these studies are limited in their context (e.g., practitioners’ organisation and knowledge or software artefacts), making it difficult to generalise the results to broader software documentation artefacts. Another source of information for researchers lies in these software artefacts themselves. Different software documentation artefacts have been evaluated, revealing a series of issues. Wen et al. [14] conducted a large-scale empirical study of code-comment inconsistencies, identifying the taxonomy of inconsistencies fixed by developers. Mirhosseini and Parnin [4] evaluated the executability of software tutorials and discussed the categories of errors that can lead to the failure of code block executions. Aghajani et al. [5] conducted a qualitative study on software documentation issues from Stack Overflow, GitHub issues, GitHub pull requests, and mailing lists, identifying both content and process issues.
@@
Among the software documentation artefacts, README files play an important role in introducing OSS repositories. They are relatively easy to access, with a rich content editing history. Despite their significance, README files have received relatively less attention in research, and most existing work focuses on more abstract perspectives, such as the structure and content topics. Prana et al. [1] conducted a qualitative study involving manual annotation of 4,226 README file sections to categorise README content into different purposes. Liu et al. [15] studied the patterns adopted for README files by GitHub Java repositories and investigated the association between repository popularity and README file structures. The study identified 32 clusters representing common README file structures, and the findings suggested that repositories adhering to GitHub guidelines for README files tend to get more stars. Meanwhile, Wang et al. [16] investigated the factors that influence the popularity of GitHub repositories from the README file angle, and identified significant difference between popular and non-popular repositories in README file related featrues.
@@
2.2 Automated Tools for Software Documentation
@@
Previous research focuses on supporting various software documentation artefacts with non-documentation artefact, such as source code. To begin with, code summarisation is an actively evolving field, with techniques developed for generating code comments [17, 18, 19, 20], commit messages [21, 22, 23], and release notes [24, 25]. Technical debt, including code comment inconsistency detection [26, 27], and self-admitted technical debt [28, 29], is also investigated to facilitate software maintenance.
Efforts have also been made to improve and augment software documentation content and navigation across this content. Treude and Robillard [30] investigated a technique for augmenting the content of API documents by incorporating information from Stack Overflow. Treude et al. [31] extracted development tasks from software documents, helping bridge the gap between software documentation and the information needs of software developers. Gao et al. [32] mined README files from GitHub and performed transfer learning to simplify complicated documentation sentences. In fact, Robillard et al. [33] have proposed a paradigm shift for automatic documentation generation based ondevelopers’ queries, given their contexts.
However, to facilitate more effective automated documentation-writing tools, it is essential to have a deeper understanding not only of the software documentation artefacts, but also of the patterns of behaviours associated with their updates. In this work, we conducted qualitative analysis of installation-related documentation update behaviours, with the aim of gaining a deep understanding and providing motivation for developing automated software tools for documentation writing.
2.3 OSS Barriers and Project Installation
@@
Moreover, another study [1] pointed out that the major content category of README files is “How”, which includes instructions from setting up the environment to installation and running the project. Therefore, README files play a vital role in guiding newcomers and software users for installation-related topics.
In contrast to previous research focusing on the human perspective of OSS barriers for newcomers, this work synthesises empirical evidence on software documentation activities, specifically on parts related to the installation, based on evidence derived from a large-scale dataset that we mined. Such empirical evidence will be useful to mitigate documentation-related barriers for newcomers when engaging with OSS projects.
3 Study Design and Methodology
The goal of this study is to investigate how installation-related instructions in README files are updated. We formulate a research question (RQ) to guide our study:
RQ: How are installation sections in README files modified? This RQ will reveal different modification behaviours for updating the documentation. We perform qualitative analysis of the installation-related sections within our mined data to answer this question. The answer to this RQ will enhance our understanding of common practice when updating the README files and provide recommendations and insights for practitioners and researchers.
3.1 Data Preparation
Python and Java are two widely used programming languages that span a broad range of applications, from developing easy-to-use tools to large-scale servers. Given that the focus of this paper is on installation-related content, which demands a good level of programming language proficiency, Python and Java were chosen because of their alignment with the expertise of all the authors. Therefore, we only consider GitHub repositories that use Python or Java as their main programming language.
Although the GitHub API provides access to its data, the rate limit is not enough for extensively collecting all repositories that meet our needs. Therefore, we used GH Archive, [^3] https://www.gharchive.org/ which collects GitHub data based on the Event API. Figure 1 describes our overall repository collection procedure.
@@ 3.1 Data Preparation
As our focus is on update behaviours, we started with commits rather than repository content. In addition, we want the repositories retrieved to contain a history of updating their installation-related instructions. Therefore, we looked at the “PushEvent” in the data records, and retrieved those with commit messages containing keywords that we pre-defined after tokenisation and lemmatisation. Keywords were derived from four base keywords “instruct”, “guide”, “install” and “deploy”. These terms were chosen to represent the content genres and topics that we aimed to include in our analysis. They were further combined with tags used in Zahedi et al. [37]. We expand the set of keywords by finding both the noun and verb forms of words in WordNet [38]. This procedure resulted in a total of 21,238,948 repositories. After that, we preserved only 5,335,695 repositories that are not forks and have at least one pull request in the history. Further refinement was applied, narrowing down the selection to 886,365 by considering repositories that use Python or Java as their primary programming language. As the aim of this study is to investigate modifications within the README file, we focused only on repositories with at least 500 commits [14] in order to obtain diverse update behaviours throughout different phases of the software development lifecycle. We also used 10 stars as a threshold for eliminating toy projects [39]. This further reduces the number of repositories to 28,896.
@@
There are caveats that need to be taken care of when mining software repositories on GH Archive. First, some repositories are no longer available, either because they have been deleted or made private by their owner. We found 254,236 of these repositories and ignored them in our analysis. Second, due to renaming and other factors, some repositories are redirected by the GitHub API to alternative names, potentially causing duplicate repository issues. We further eliminated these repositories, reducing the number of repositories to 13,245.
In the last step, we removed outliers that fall outside of the scope of this study. We used FastText [40] to filter out 292 repositories with non-English README content. We manually inspected the repositories included after these filtering steps and found that there were some non-software repositories in there. Non-software repositories refer to projects primarily contain content other than traditional software code. These repositories may include, but are not limited to, collections of awesome lists, educational materials, and personal websites. To exclude those systematically, we considered several factors that might indicate non-software projects, including (1) README commits to overall commits ratio, (2) number of README commits, (3) number of README commits per day, and (4) repository size. We sorted the repositories by these factors in descending order and manually labelled the top 100 records for each factor, respectively. Figure 2 illustrates how the percentage of non-software repositories varies as more projects are included. Although the number of README commits can identify the most unrelated repositories, the accuracy is merely 51% for the top 100 entries. Meanwhile, regarding README commits to overall commit ratio, a plateau on the left side of the red vertical line indicates a consistent high accuracy of identifying non-software projects, with the red line representing the percentage of 52.9% of this ratio. Therefore, we rounded the ratio to 50%, used it as a threshold, and discarded these repositories, eventually obtaining a dataset consisting of 12,908 repositories.
@@
3.2 README Commit Classification
We randomly sampled 400 repositories from the entire population, ensuring a confidence level above 95%. For the 400 repositories sampled, there are 25,209 README commits in total, which need to be classified as “relevant” or “irrelevant” to the installation-related instructions. Given the prohibitively large number for manual annotation, we employed a semi-automated heuristic to significantly reduce the manual workload for identifying the relevant README updates.

### v3

# Adapting Installation Instructions in Rapidly Evolving Software Ecosystems

@@ Abstract
README files play an important role in providing installation-related instructions to software users and are widely used in open source software systems on platforms such as GitHub. Software projects evolve rapidly alongside their dependencies in dynamic software ecosystems, requiring frequent updates to installation instructions. These instructions are crucial for users to start with a software project. Despite their significance, there is a lack of systematic understanding regarding the documentation efforts invested in README files and the triggers behind them. To fill the research gap, we conducted a qualitative study, investigating 400 GitHub repositories with 1,163 README commits that focused on updates in installation-related sections. Our research revealed six major categories of changes in the README commits, namely pre-installation instructions, installation instructions, post-installation instructions, help information updates, document presentation, and external resource management. We further provide detailed insights into modification behaviours and offer examples of these updates. We also studied the triggers for the documentation updates, which led to three categories including errors in the previous documentation, changes in the codebase, and need for documentation improvement. Based on our findings, we proposed a README template tailored to cover the installation-related sections for documentation maintainers to reference when updating documents. We further validated this template by conducting an online survey and a pull request study, identifying that documentation readers find the augmented documents based on our template to be generally of better quality, and documentation maintainers find it useful. We further provide recommendations to practitioners for maintaining their README files, as well as motivations for future research directions. These recommendations encompass completeness, correctness and up-to-dateness, and information presentation considerations. The proposed research directions include the development of automated tools, in particular for documentation updates, and conducting empirical studies to enhance comprehension of the needs of documentation users.
@@
Software Documentation, README Files, Qualitative Analysis, Open Source Software, Installation Instructions, Software Ecosystem
I Introduction
@@
Good software documentation serves as a vital source of information for different stakeholders involved in software systems. In particular, README files are among the first software documents that users, such as other developers and end-users, and newcomers encounter when accessing an Open Source Software (OSS) project, leaving them with a crucial first impression and providing them with an introductory overview of the project [1, 2]. At the time of this study, GitHub, the most popular open-source code platform, hosts more than 372 million repositories, [^1] http://web.archive.org/web/20230424004513/https://en.wikipedia.org/wiki/GitHub underscoring the high demand for maintaining high-quality README files.
@@
One of the most common types of content in GitHub README files is related to basic instructions (e.g., project installation guide) [1]. A well-written “How” section could alleviate the effort for developers and users, especially newcomers, to get hands on with the projects. In fact, overcoming the challenge of setting up an environment is one of the hurdles for newcomers to contribute to OSS projects [3].
@@
Software projects evolve rapidly alongside their dependencies in dynamic software ecosystems [4], requiring frequent updates to installation instructions. Therefore, README files need to be carefully maintained; otherwise, various software documentation issues may arise. For example, software tutorials that focus on installing software tools, configuring virtual computing environments, and other related continuous deployment processes have been found to be not executable in more than half of the cases investigated by related work [5]. Furthermore, previous empirical studies identified inappropriate installation instructions, outdated content, and content clarity as frequently appearing documentation issues [6, 7]. These factors collectively constitute the impediments for both developers and project users to use software repositories, while empirical evidence on the current documentation practices and their triggers could pave the way for mitigating these issues from source.
@@
Although previous studies have pinpointed software documentation issues, particularly in installation-related topics [6, 7, 8], their analyses were broad in scope, with the aim of synthesising different aspects of documentation issues from surveys or software documents. In contrast, documentation updates provide a new perspective consisting of the modifications and two states: before and after the modification, while triggers offer the motivations that drive documentation updates in the rapidly evolving software ecosystem. In fact, increased documentation activities and efforts have demonstrated their efficacy in improving the popularity of and encouraging participation in GitHub projects [9]. Considering the documentation issues and the benefits of investing efforts to maintain high-quality documentation, knowledge of the practices of the README file efforts and the triggers behind them becomes essential. This knowledge reveals documentation activities in the dynamic software ecosystems. It not only provides practitioners with recommendations but also provides researchers with insights for the development of automated documentation tools. However, empirical evidence to support practices related to software documentation efforts remains scarce.
@@
In this work, we conducted an empirical study to qualitatively analyse README file modification activities in installation-related topics to fill the research gap. We mined GitHub repositories from GHArchive that were active between 2015-01-01 and 2023-02-28, [^2] https://www.gharchive.org/ applying stringent filters on different features for the repositories, resulting in 12,908 projects. We examined a statistically significant sample of 1,168 README file commits within 400 randomly selected repositories and adopted qualitative analysis to synthesise the types of README update tasks in the modified sections.
@@
Based on our analysis, we synthesised a comprehensive taxonomy of 189 update behaviours, which are related to the categories of: (1) pre-installation instructions, (2) installation instructions, (3) post-installation instructions, (4) help information, (5) document presentation, and (6) external resource management. We then devised a process to identify the triggers for documentation updates. Our analysis leads to three categories of documentation update triggers: (1) errors in the previous documentation, (2) changes in the codebase, and (3) need for documentation improvement. After that, we proposed a README template tailored to cover the installation-related sections for documentation maintainers to reference when updating documents. The template validation process indicates that readers find README files that apply our template to be of generally increased quality, and the documentation maintainers also find the process of applying the template useful. This serves as a fundamental step towards automated documentation updates. We further discussed their implications for documentation maintainers and software engineering researchers based on our discoveries from perspectives of completeness, correctness and up-to-dateness, and information presentation considerations.
Our key contributions are as follows: (1) A large-scale repository dataset, which can be used for future research. (2) Formalisation of a detailed list of software development tasks related to README file content. (3) A comprehensive taxonomy for the types of installation-related documentation modifications for OSS projects. (4) A README template validated by documentation readers to be of better quality compared to original README files. (5) A categorisation of triggers for documentation updates (6) Actionable recommendations for OSS documentation maintainers and researchers, in particular the automated documentation updates direction.
The rest of the paper is structured as follows: Section 2 covers the related work; Section 3 introduces our study design and methodology adopted to conduct our study; Section 4 presents the results of our qualitative analysis for the installation instruction taxonomy; Section 5 presents the results for triggers for documentation updates; Section 6 provides the README template from our study and its validation process with an online survey; Section 7 discusses the implications of this study to practitioners and researchers; Section 8 covers the threats to the validity of our study; Section 9 provides the conclusion. Our replication package is at https://github.com/Haoyu-Gao/README_Empirical_Study, with the dataset in the root folder.
II Related Work
II-A Empirical Studies on Software Documentation
@@
One source of information for researchers is directly from software engineering practitioners. Several studies have investigated the impact of software documentation. Kajiko-Mattson [10] studied the practices of software documentation in Swedish organisations, identifying rudimentary documentation requirements relevant within corrective maintenance. Garousi et al. [11] conducted an industrial case study on the usage and usefulness of different documentation artefacts in the software development life cycle. Their findings indicated variations in documentation usage based on information needs, and the recommendations derived from their study have been integrated into the company’s practices where the experiment took place. Software documentation issues are also investigated. Uddin and Robillard [12] conducted a large-scale survey on API documentation failure factors among IBM practitioners, and identified the most pressing issues related to content, as opposed to presentation. Aghajani et al. [7] presented the views of practitioners on various issues of software documentation, along with their proposed solutions to address these issues. Moreover, the quality of software documents is another perspective that has been studied. Treude et al. [13] recruited technical editors and synthesised a framework to assess the quality of the software documentation from the respondents. Their work provides visions for a potential unified quality framework for assessing software documentation.
@@
However, these studies are limited in their context (e.g., practitioners’ organisation and knowledge or software artefacts), making it difficult to generalise the results to broader software documentation artefacts. Another source of information for researchers lies in these software artefacts themselves. Different software documentation artefacts have been evaluated, revealing a series of issues. Wen et al. [14] conducted a large-scale empirical study of code-comment inconsistencies, identifying the taxonomy of inconsistencies fixed by developers. Mirhosseini and Parnin [5] evaluated the executability of software tutorials and discussed the categories of errors that can lead to the failure of code block executions. Aghajani et al. [6] conducted a qualitative study on software documentation issues from Stack Overflow, GitHub issues, GitHub pull requests, and mailing lists, identifying both content and process issues. Gao et al. [15] studied machine learning documentation to highlight the insufficient effort in documenting ethical considerations.
@@
Among the software documentation artefacts, README files play an important role in introducing OSS repositories. They are relatively easy to access, with a rich content editing history. Despite their significance, README files have received relatively less attention in research, and most existing work focuses on more abstract perspectives, such as structure and content topics. Prana et al. [1] conducted a qualitative study involving manual annotation of 4,226 README file sections to categorise README content into different purposes. Liu et al. [16] studied the patterns adopted for README files by GitHub Java repositories and investigated the association between repository popularity and README file structures. The study identified 32 clusters representing common README file structures, and the findings suggested that repositories adhering to GitHub guidelines for README files tend to get more stars. Meanwhile, Wang et al. [17] investigated the factors that influence the popularity of GitHub repositories from the README file angle, and identified a significant difference between popular and non-popular repositories in README file-related features.
@@
II-B Automated Tools for Software Documentation
@@
Previous research focuses on supporting various software documentation artefacts with non-documentation artefacts, such as source code. To begin with, code summarisation is an actively evolving field, with techniques developed for generating code comments [18, 19, 20, 21], commit messages [22, 23, 24], code reviews [25, 26] and release notes [27, 28]. Technical debt, including code comment inconsistency detection [29, 30], and self-admitted technical debt [31, 32], is also investigated to facilitate software maintenance.
Efforts have also been made to improve and augment software documentation content and navigation across this content. Treude and Robillard [33] investigated a technique for augmenting the content of API documents by incorporating information from Stack Overflow. Treude et al. [34] extracted development tasks from software documents, bridging the gap between software documentation and the information needs of software developers. Gao et al. [35] mined README files from GitHub and performed transfer learning to simplify complicated documentation sentences. In fact, Robillard et al. [36] have proposed a paradigm shift for automatic documentation generation based on developers’ queries, given their contexts.
However, to facilitate more effective automated documentation-writing tools, it is essential to have a deeper understanding not only of the software documentation artefacts but also of the patterns of behaviours associated with their updates. In this work, we conducted a qualitative analysis of installation-related documentation update behaviours along with the triggers behind them, with the aim of gaining a deep understanding and providing motivation for developing automated software tools for documentation writing.
II-C OSS Barriers and Project Installation
@@
Moreover, another study [1] pointed out that the major content category of README files is “How”, which includes instructions from setting up the environment to installation and running the project. Therefore, README files play a vital role in guiding newcomers and software users with installation-related topics.
In contrast to previous research on the human perspective of OSS barriers for newcomers, this work synthesises empirical evidence on software documentation activities, specifically on parts related to the installation, based on evidence derived from a large-scale dataset that we mined. Such empirical evidence will be useful to mitigate documentation-related barriers for newcomers when engaging with OSS projects.
III Study Design and Methodology
The goal of this study is to investigate how installation-related instructions in README files are updated. We formulate three research questions (RQs) to guide our study:
RQ1: How are installation sections in README files modified? This RQ will reveal different modification behaviours for updating the documentation. We studied the documentation modification behaviour rather than the most recent version of README files because the modification behaviours not only provide the documentation structure and contents but also offer more insights into the documentation maintenance efforts on OSS repositories. By understanding the documentation maintenance practice, we could provide implications to practitioners and researchers for better maintaining the README files. We performed a qualitative analysis of the installation-related sections within our mined data to answer this question. One of the target outcomes of this RQ is a template encompassing all essential components to facilitate better documentation activities for the maintainers.
RQ2: What are the triggers for the README updates in the installation sections? This RQ investigates the reasons that trigger documentation updates in the software system, providing insights into what drives documentation updates in a rapidly evolving software ecosystem. The triggers and update behaviours collectively provide strong empirical evidence for automatic documentation updates as a future direction.
RQ3: Could a template for installation instructions be devised to improve documentation quality? This research question aims to validate our proposed README template, which is the fundamental first step towards automated documentation updates. We conducted a survey asking participants to evaluate the perceived documentation quality before and after applying the README template, based on an established documentation assessment framework [13]. We further conducted a pull request study, similar to Malaquias et al. [40] to evaluate the practitioners’ perspective on our template.
III-A Data Preparation
In this paper, we only consider GitHub repositories that use Python or Java as their main programming language. Python and Java are two widely used programming languages that span a broad range of applications, from developing easy-to-use tools to large-scale servers. In fact, Python and Java consistently rank among the top three most used programming languages on GitHub, [^3] https://octoverse.github.com/2022/top-programming-language following JavaScript.
Although the GitHub API provides access to its data, the rate limit is not enough for collecting all repositories that meet our needs. Therefore, we used GH Archive, [^4] https://www.gharchive.org/ which collects GitHub data based on the Event API. Figure 1 describes our overall repository collection procedure.
@@ III-A Data Preparation
As our focus is on update behaviours, we started with commits rather than repository content. In addition, we want the repositories retrieved to contain a history of updating their installation-related instructions. Therefore, we looked at the “PushEvent” in the data records, and retrieved those with commit messages containing keywords that we pre-defined after tokenisation and lemmatisation. Keywords were derived from four base keywords “instruct”, “guide”, “install” and “deploy”. These terms were chosen to represent the content genres and topics that we aimed to include in our analysis. They were further combined with tags used in Zahedi et al. [41]. We expand the set of keywords by finding both the noun and verb forms of words in WordNet [42]. This procedure resulted in a total of 21,238,948 repositories. After that, we preserved only 5,335,695 repositories that are not forks and have at least one pull request in the history. This filtering step ensures that the selected repositories demonstrate development activity, rather than being solely for archival purposes or mirrors of other open-source platforms such as BitBucket or GitLab [43]. Further refinement narrows down the selection to 886,365 by considering repositories that use Python or Java as their primary programming language. As the aim of this study is to investigate modifications within the README file, we focus only on repositories with at least 500 commits [14] to obtain repositories that went through code changes in different phases of the software development lifecycle. In this way, the README updates ideally capture a broader range of changes. Sampling from the entire population would give higher weight to repositories with fewer commits, potentially reducing diversity in the types of change reflected in README updates. We also used 10 stars as a threshold for eliminating toy projects [44]. This further reduces the number of repositories to 28,896.
@@
There are caveats that need to be taken care of when mining software repositories on GH Archive. First, some repositories are no longer available, either because they have been deleted or made private by their owners. After removing these non-available repositories, we retain 14,485 repositories. Second, due to renaming and other factors, some repositories are redirected by the GitHub API to alternative names, potentially causing duplicate repository issues. We further eliminated these repositories, reducing the number of repositories to 13,245.
In the last step, we removed outliers that fall outside of the scope of this study. We used FastText [45] to filter out 292 repositories with non-English README content. We manually inspected the repositories included after these filtering steps and found that there were some non-software repositories in there. Non-software repositories refer to projects primarily containing content other than traditional software code. These repositories may include, but are not limited to, collections of awesome lists, educational materials, and personal websites. To exclude those systematically, we considered several factors that might indicate non-software projects, including (1) README commits to overall commits ratio, (2) number of README commits, (3) number of README commits per day, and (4) repository size. We sorted the repositories by these factors in descending order and manually labelled the top 100 records for each factor, respectively. Figure 2 illustrates how the percentage of non-software repositories varies as more projects are included. Although the number of README commits can identify the most unrelated repositories, the accuracy is merely 51% for the top 100 entries. Meanwhile, regarding README commits to overall commit ratio, a plateau on the left side of the red vertical line indicates a consistently high accuracy of identifying non-software projects, with the red line representing the percentage of 52.9% of this ratio. Therefore, we rounded the ratio to 50%, used it as a threshold, and discarded these repositories, eventually obtaining a dataset consisting of 12,908 repositories.
@@
III-B README Commit Classification
We randomly sampled 400 repositories from the entire population, ensuring a confidence level above 95% with a 5% margin of error. For the 400 repositories sampled, there are 25,209 README commits in total, which need to be classified as “relevant” or “irrelevant” to the installation-related instructions. Given the prohibitively large number for manual annotation, we employed a semi-automated heuristic to significantly reduce the manual workload for identifying the relevant README updates.